# Imprecise Markov Semigroups and their Ergodicity

### v4

@@ 2. Imprecise Markov Semigroups
Let us add two remarks. First, the reader may ask themselves why we went to such great lengths to ensure that $\mathbf{Q}$ possesses the least and greatest elements $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$, respectively. Informally, the reason is that studying their limiting behavior (that is, studying the limits $\lim_{t\rightarrow\infty}\underline{P}_{t}\tilde{f}$ and $\lim_{t\rightarrow\infty}\overline{P}_{t}\tilde{f}$) will inform us on the limiting behavior of all the elements of $\mathbf{Q}$ evaluated at $\tilde{f}$. This intuition is made formal in Corollaries 8.1 and 10.1. In addition, since $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$ are (precise) Markov semigroups, we can use techniques developed e.g. in Bakry et al. (2014) to carry out such a study.
[…]
It is also worth mentioning that we focus our attention on orders (total and partial) associated with a particular function $\tilde{f}$ because, given an order of the type $\mathbf{P}\preceq\mathbf{P^{\prime}}\iff P_{t}f\leq P^{\prime}_{t}f$, for all $t\geq 0$ and all $f\in B(E)$, there exists at least one set $\mathbf{Q}$ that satisfies (1), but which does not possess a least element $\underline{\mathbf{P}}$, according to such an order (Krak et al., 2017, Example 6.2).
@@
Let us pause here to expand on the last point. Working with Imprecise Markov Semigroups (IMSGs) allows us to capture the ambiguity faced by the agent in assessing both the invariant measure and the transition probability of a Markov process. By studying IMSGs, we take a different perspective to imprecise Markov processes than the ones usually employed in the literature introduced in Section 1. Indeed, existing works model invariant measure ambiguity via an invariant Choquet capacity – see e.g. de Cooman et al. (2009) – and transition probability ambiguity via sets of transition probabilities or sets of transition rate matrices, usually assumed convex to allow maximization and minimization operators to be computed in terms of linear programming optimization problems – see e.g. Krak et al. (2017, 2019). In this paper, instead, we capture both types of ambiguities via the imprecise Markov semigroup $\mathbf{Q}$. From Definition 7, the latter is a collection of (precise) Markov semigroups $\mathbf{P}=(P_{t})_{t\geq 0}$, each associated with a Markov process $(X_{t})_{t\geq 0}$. In turn, every such Markov semigroup $\mathbf{P}$ encompasses a plausible invariant measure and a plausible transition probability. The scholar, then, can express their ambiguity around both, by letting the Markov processes $(X_{t})_{t\geq 0}$ associated with the elements $\mathbf{P}$ of $\mathbf{Q}$ have different invariant measures and transition probabilities.
@@
As we shall see in Sections 3 and 4, taking an (imprecise Markov) semigroup route allows us to explicitly account for the geometry of the state space $E$ when we study the ergodic behavior of $\overline{\mathbf{P}}$ and $\underline{\mathbf{P}}$ (a perspective that is new to the literature), and to work in continuous time (whereas the majority of the results in the literature pertain the discrete time case). The price we pay for these gains is the need for a compact set of semigroups of linear operators. This is because, by Lemma 6, $\overline{\mathbf{P}},\underline{\mathbf{P}}\in\mathbf{Q}$, and so we can use techniques developed for (precise) Markov semigroups to study the ergodic behavior of $\overline{\mathbf{P}}$ and $\underline{\mathbf{P}}$, and in turn of all the elements of $\mathbf{Q}$. In the future, we plan to extend our approach involving the geometry of the state space $E$ to the frameworks introduced e.g. in Criens and Niemann (2024); Denk et al. (2018, 2020); Erreygers (2023); Feng and Zhao (2021); Kühn (2021); Nendel (2021), [^7] In those papers, the curvature of the state space $E$ is not considered by the authors when deriving their results. where nonlinear Markov semigroups (and in particular Markov semigroups associated with nonlinear Markov processes) are studied. This would allow us to retain the advantages of the semigroup approach, but without the restrictions coming from enforcing compactness of the imprecise Markov semigroup $\mathbf{Q}$, and linearity of its elements.
3. Ergodicity in the Riemannian Manifold Case
@@
Loosely, a complete Riemannian manifold $M$ can be thought of as a “smooth” subset of $\mathbb{R}^{n}$, equipped with a complete metric that depends in some sense on the elements $x\in M$. A visual representation of a Riemannian manifold in $\mathbb{R}^{3}$ is given in Figure 3. In this context, the carré du champ and the iterated carré du champ operators associated with a Markov semigroup are easily computed (see B.1.1).
@@
Let $E$ be $\mathbb{R}^{n}$ or a smooth complete connected manifold. Let $\mathcal{A}_{0}$ be the class of smooth ($\mathcal{C}^{\infty}$) compactly supported functionals $f$ on EE. [^9] Recall that a smooth function on a closed and bounded set (and hence, in this framework, compact by Heine-Borel) is itself bounded. Fix any $\tilde{f}\in B(E)$, and assume that we can define the partial order $\preceq_{\tilde{f}}^{\text{part}}$ or the total order $\preceq_{\tilde{f}}^{\text{tot}}$ that we introduced in Section 2. Let $\mathbf{Q}$ be an imprecise Markov semigroup. Denote by $\overline{\mathbf{P}}$ and $\underline{\mathbf{P}}$ the greatest and least elements of $\mathbf{Q}$, [^10] With respect to $\preceq^{\text{part}}_{\tilde{f}}$ or $\preceq^{\text{tot}}_{\tilde{f}}$. respectively, with invariant measures $\overline{\mu}$ and $\underline{\mu}$, and associated infinitesimal generators $\overline{L}$ and $\underline{L}$, respectively. Suppose that the following three conditions hold
@@
If $E$ is a complete Riemannian manifold, the Bakry-Émery curvature condition $CD(\rho_{\underline{L}},\infty)$ in Theorem 8.(ii) is related to the notion of Ricci tensor at every point of $E$, and in turn to the Ricci curvature of the manifold $E$ (for more, see B.1.1). [^12] A similar argument holds for the infinitesimal generator $\overline{L}$ associated with Markov semigroup $\overline{\mathbf{P}}$. We see, then, how the geometry of the state space $E$ influences the ergodic behavior of an imprecise Markov semigroup.
Figure 3. In this figure, $M$ is a Riemannian manifold in $\mathbb{R}^{3}$, $x\in M$ is an element of the manifold, and $T_{x}M$ is the tangent space of $M$ at $x$ (the set of all tangent vectors to all smooth curves on $M$ passing through $x$). $\gamma(t)$ is a $\mathcal{C}^{1}$ curve passing through $x$, and $v\equiv v(x)=(v^{i}(x))_{i\in\{1,2,3\}}$ is a vector field.
@@
We begin by pointing out that, by Lemma 6, $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$ are (precise) Markov semigroups, and so the infinitesimal generators $\underline{L}$ and $\overline{L}$ are well-defined.
@@
Following van Handel (2016, Problem 2.7), taking the limit as $t\rightarrow\infty$ on both sides of (5) yields the following Poincaré inequality
@@
In addition, recall that for the Dirichlet form the Dirichlet form $\overline{\mathcal{E}}(f)$ we have that
@@
Combining (7) with (6), we have that if function $f$ is such that $\overline{L}f=0$, then $\overline{\mathcal{E}}(f)=0$. In turn, this implies that $\text{Var}_{\overline{\mu}}(f)=0$, and so the function $f$ must be constant $\overline{\mu}$-a.e. Notice now that, given that $\overline{L}$ is an elliptic operator, $\mathcal{A}_{0}$ is dense in $\mathcal{D}(\overline{L})$ in the topology induced by the infinitesimal generator domain norm $\|\cdot\|_{\mathcal{D}(\overline{L})}$. [^13] This norm is recalled in Definition 18.(DMT9) in B.2. Thanks to this, (3)-(7) hold for all $f\in\mathcal{D}(\overline{L})$ by an approximation argument. This implies that every $f\in\mathcal{D}(\overline{L})$ such that $\overline{L}f=0$ is constant, and so that $\overline{L}$ is ergodic according to Bakry et al. (2014, Definition 3.1.11). Then, since $\overline{\mu}$ is finite, by Bakry et al. (2014, Proposition 3.1.13) we have that
@@
Examples of semigroups for which the conditions (i)-(iii) of Theorem 8 are met are the Ornstein-Uhlenbeck and Laguerre semigroups (whose Markov generators satisfy the Bakry-Émery curvature condition with $\rho>0$; in particular, $\rho=1$ for the former, and $\rho=1/2$ for the latter), and the Brownian (Euclidean) semigroup (whose Markov generator satisfies the $CD(\rho,\infty)$ condition with $\rho=0$). In addition, an example of two Markov semigroups having the same invariant measure (as required for (2) to hold) are the Brownian motion with quadratic potential, and the Ornstein-Uhlenbeck semigroup with volatility parameter $\sigma=\sqrt{2}$. Indeed, for both, the invariant measure is a Gaussian with mean 0, and variance the inverse of the rate of mean reversion.
The following is an important corollary that subsumes the ergodic behavior of the elements of $\mathbf{Q}$ when applied to $\tilde{f}\in B(E)$.
Corollary 8.1 (Ergodicity of Imprecise Markov Semigroup $\mathbf{Q}$, $E$ Euclidean space or Riemannian manifold).
@@
Let us add some interesting comments on Corollary 8.1. In (10), we see how, if the initial distribution – i.e. the distribution of $X_{0}$ – is known and equal to $\mu$, then the ambiguity that the agent faces around the correct transition probability of the process from $X_{0}$ to $X_{t}$, vanishes as $t\rightarrow\infty$. We recover the initial (reversible invariant) distribution $\mu$, according to which we compute the expectations. As the classical ergodic adage goes, “in the limit, time average ($\lim_{t\rightarrow\infty}\underline{P}_{t}\tilde{f}=\lim_{t\rightarrow\infty}\overline{P}_{t}\tilde{f}$) equals space average ($\int_{E}\tilde{f}\text{d}\mu$)”. We say that an imprecise Markov semigroup $\mathbf{Q}$ is $\tilde{f}-$lower ergodic if it satisfies (8), $\tilde{f}-$upper ergodic if it satisfies (9), and $\tilde{f}-$ergodic if it satisfies (10).
As we pointed out in Section 2, there are two main differences with respect to the existing literatures on imprecise Markov processes and ergodicity for imprecise probabilities. We are the first to explicitly exploit the geometry of the state space $E$ via the Bakry-Émery curvature condition to derive our results. In addition, in Theorem 8 and Corollary 8.1 we represent the scholar’s ambiguity about the invariant measure and the transition probability via an imprecise Markov semigroup $\mathbf{Q}$. We briefly mention that our approach differs slightly from the imprecise ergodic literature (see e.g. Caprio and Gong (2023); Caprio and Mukherjee (2023); Cerreia-Vioglio et al. (2015)), where ambiguity around distribution $\mu$ is typically represented via a Choquet approach involving a superadditive Choquet capacity (Choquet, 1954).
We conclude this section with a corollary that further informs us on the geometrical properties of $E$ in the ergodic regime of $\mathbf{Q}$, when $E$ is a complete Riemannian manifold. Let us denote by $E^{n}_{K}$ the complete $n$-dimensional simply connected space of constant sectional curvature $K\in\mathbb{R}$, [^14] The Ricci curvature of $E^{n}_{K}$ is constant and equal to (n1)K(n-1)K. and by $B(p,r)$ the ball of radius $r$ at point $p$, defined with respect to the Riemannian distance function.
Corollary 8.2 (Bishop–Gromov Inequality).
@@
Given our assumptions, we know that $CD(\rho_{\underline{L}},\infty)$ and $CD(\rho_{\overline{L}},\infty)$ both hold. Let us define $\rho_{\star}\coloneqq\min\{\rho_{\underline{L}},\rho_{\overline{L}}\}$, and put $\rho_{\star}=(n-1)K$, so that $K=\rho_{\star}/(n-1)$. The Bakry-Émery curvature assumptions imply that the Ricci tensor at every point of $E$ is bounded from below by $\rho_{\star}$. [^15] This is discussed in B.1.1. In turn, we have that the Ricci curvature Ric of $E$ is lower bounded by (n1)K(n-1)K, that is, $\text{Ric}\geq(n-1)K$. The result then follows from Bishop (1963). ∎
@@
4. Ergodicity in the General Case
In this section, our goal is to extend Theorem 8 and Corollary 8.1 to the case where $(E,\mathcal{F})$ is an arbitrary good measurable space. To do so, we need to introduce the concept of an Imprecise Diffusion Markov Tuple, the imprecise version of a diffusion Markov triple. The latter is recalled in B.2.

### v5

@@ 2 Imprecise Markov Semigroups
At this point, the reader may ask themselves why we went to such great lengths to ensure that $\mathbf{Q}$ possesses the least and greatest elements $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$, respectively. Informally, the reason is that studying their limiting behavior (that is, studying the limits $\lim_{t\rightarrow\infty}\underline{P}_{t}\tilde{f}$ and $\lim_{t\rightarrow\infty}\overline{P}_{t}\tilde{f}$) will inform us on the limiting behavior of all the elements of $\mathbf{Q}$ evaluated at $\tilde{f}$. This intuition is made formal in Corollaries 10.1 and 12.1. In addition, since $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$ are (precise) Markov semigroups, we can use techniques developed in the Markov diffusion operators literature (Bakry et al., 2014) to carry out such a study.
Notice also that the framework of Definition 9 is deliberately non-switching. That is, we model ambiguity by a set of precise Markov semigroups $\mathbf{Q}$, but we do not allow the decision-maker to concatenate two different models over successive time intervals. Concretely, given $\mathbf{P},\mathbf{P}^{\prime}\in\mathbf{Q}$, we do not assume that there exists $\mathbf{P}^{\prime\prime}\in\mathbf{Q}$ such that
$P^{\prime\prime}_{t+s}=P^{\prime}_{t}\circ P_{s},\quad\text{ for some }t,s\geq 0,$
nor do we interpret $P^{\prime}_{t}$ as acting on the random output $P_{s}\tilde{f}$ produced under a different model $\mathbf{P}$. Allowing such concatenations corresponds to admitting switching controls, and typically leads to a time-consistent upper (or lower) expectation that is no longer representable by a single precise semigroup: The resulting one-step kernels need not be dominated by a fixed probability kernel and may instead be described by nonlinear transition operators – e.g. capacities in the sense of Choquet (1954).
In Sections 3 and 4 we therefore restrict attention to ambiguity represented by a collection of precise semigroups without switching. By contrast, in Section 5 we explicitly allow switching and construct the associated (Nisio) nonlinear semigroups, which lets us dispense with compactness while retaining robust long-term behavior.
Furthermore, it is worth mentioning that in Sections 3 and 4 we focus our attention on orders (total and partial) associated with a particular function $\tilde{f}$. Indeed, if one were to consider the “global” comparison
$\mathbf{P}\preceq\mathbf{P}^{\prime}\quad\iff\quad P_{t}f\leq P_{t}^{\prime}f,\text{ for all }t\geq 0\text{ and all }f\in B(E),$
then this relation is in fact trivial: Applying it to f-f yields $P_{t}f\geq P_{t}^{\prime}f$, hence $P_{t}f=P_{t}^{\prime}f$ for all $t\geq 0$ and $f\in B(E)$. In other words, such an order collapses to equality and cannot distinguish distinct semigroups. Therefore, to obtain a nontrivial comparison and meaningful extremal elements, one must restrict the function class, which we do here by working with orders induced by a fixed $\tilde{f}$. We note that even under a restricted order, there are families $\mathbf{Q}$ satisfying (1) that still fail to possess a least element $\underline{\mathbf{P}}$ with respect to that order, see Krak et al. (2017, Example 6.2).
@@
In the imprecise probability literature, ambiguity about invariant behavior is often encoded directly at the level of stationary objects, for instance through an invariant Choquet capacity (de Cooman et al., 2009). Ambiguity about the dynamics is instead typically modeled by sets of transition probabilities (in discrete time) or sets of transition rate matrices (in continuous time), frequently assumed convex so that worst/best-case evaluations can be computed via linear-programming-type procedures (Krak et al., 2017, 2019).
A Markov semigroup does not, by itself, specify an initial distribution: It encodes only the transition mechanism $(P_{t})_{t\geq 0}$, while an initial law $\nu\in\Delta_{E}$ must be supplied separately to determine the full process law; Here $\Delta_{E}$ denotes the space of countably additive probability measures on $E$. In this paper, we represent ambiguity about the dynamics by an imprecise Markov semigroup $\mathbf{Q}$, i.e. a family of precise semigroups $\mathbf{P}=(P_{t})_{t\geq 0}$. Each $\mathbf{P}\in\mathbf{Q}$ determines a candidate collection of transition probabilities and (when it exists) an associated invariant measure $\mu\equiv\mu_{\mathbf{P}}$. Hence $\mathbf{Q}$ captures ambiguity both about transition probabilities and about invariant/long-run behavior, by allowing these objects to vary across $\mathbf{P}\in\mathbf{Q}$. If one also wishes to model ambiguity about the initial distribution, this can be done exogenously by specifying a set of priors $\mathsf{N}\subseteq\Delta_{E}$ and considering $\{\nu P_{t}:\nu\in\mathsf{N},\mathbf{P}\in\mathbf{Q}\}$; Our results primarily focus on the robust long-term behavior induced by uncertainty in the semigroup itself.
@@
As we shall see in Sections 3, 4, and 5, taking an (imprecise Markov) semigroup route allows us to explicitly account for the geometry of the state space $E$ when we study the ergodic behavior of $\overline{\mathbf{P}}$ and $\underline{\mathbf{P}}$ (a perspective that is new to the literature), and to work in continuous time (whereas the majority of the results in the literature pertain the discrete time case). The price we pay in Sections 3 and 4 for these gains is the need for compactness of the equivalence class space under our chosen preorder. This is because, as a consequence of Lemma 7, $\overline{\mathbf{P}},\underline{\mathbf{P}}\in\mathbf{Q}$, and so we can use techniques developed for (precise) Markov semigroups to study the ergodic behavior of $\overline{\mathbf{P}}$ and $\underline{\mathbf{P}}$, and in turn of all the elements of $\mathbf{Q}$. In Section 5, instead, we allow $\mathbf{Q}$ to be non-compact, thus making our approach closer to the frameworks introduced e.g. in Criens and Niemann (2024); Denk et al. (2018, 2020); Erreygers (2023); Feng and Zhao (2021); Kühn (2021); Nendel (2021), [^4] In those papers, the curvature of the state space $E$ is not considered by the authors when deriving their results. where nonlinear Markov semigroups (and in particular Markov semigroups associated with nonlinear Markov processes) are studied.
3 Ergodicity in the Riemannian Manifold Case – Compact $\mathbf{Q}$
@@
Loosely, a complete Riemannian manifold $M$ can be thought of as a “smooth” subset of $\mathbb{R}^{n}$, equipped with a complete metric that depends in some sense on the elements $x\in M$. A visual representation of a Riemannian manifold in $\mathbb{R}^{3}$ is given in Figure 3. In this context, the carré du champ and the iterated carré du champ operators associated with a Markov semigroup are easily computed (see Appendix C.1.1). Recall from Remark 3 that, in the following theorem, we have $\underline{\mathbf{P}}\equiv\underline{\mathbf{P}}_{\tilde{f}}$, a fixed chosen representative from $\underline{[\mathbf{P}]}_{\tilde{f}}$, and similarly for $\overline{\mathbf{P}}$.
@@
Let $E$ be $\mathbb{R}^{n}$ or a smooth complete connected manifold. Let $\mathcal{A}_{0}$ be the class of smooth ($\mathcal{C}^{\infty}$) compactly supported functionals $f$ on EE. [^6] Recall that a smooth function on a closed and bounded set (and hence, in this framework, compact by Heine-Borel, or Hopf-Rinow for Riemannian manifold) is itself bounded. Fix any $\tilde{f}\in B(E)$, and assume that we can define the partial order $\preceq_{\tilde{f}}^{\text{part}}$ or the total order $\preceq_{\tilde{f}}^{\text{tot}}$ that we introduced in Section 2. Let $\mathbf{Q}$ be a compact imprecise Markov semigroup. Denote by $\overline{\mathbf{P}}$ and $\underline{\mathbf{P}}$ the greatest and least elements of $\mathbf{Q}$, [^7] With respect to $\preceq^{\text{part}}_{\tilde{f}}$ or $\preceq^{\text{tot}}_{\tilde{f}}$. respectively, with invariant measures $\overline{\mu}$ and $\underline{\mu}$, and associated infinitesimal generators $\overline{L}$ and $\underline{L}$, respectively. Suppose that the following three conditions hold,
@@
If $E$ is a complete Riemannian manifold, the Bakry-Émery curvature condition $CD(\rho_{\underline{L}},\infty)$ in Theorem 10.(ii) is related to the notion of Ricci tensor at every point of $E$, and in turn to the Ricci curvature of the manifold $E$ (for more, we refer the reader to Appendix C.1.1). [^9] A similar argument holds for the infinitesimal generator $\overline{L}$ associated with Markov semigroup $\overline{\mathbf{P}}$. We see, then, how the geometry of the state space $E$ influences the ergodic behavior of an imprecise Markov semigroup.
Figure 3: In this figure, $M$ is a Riemannian manifold in $\mathbb{R}^{3}$, $x\in M$ is an element of the manifold, and $T_{x}M$ is the tangent space of $M$ at $x$ (the set of all tangent vectors to all smooth curves on $M$ passing through $x$). $\gamma(t)$ is a $\mathcal{C}^{1}$ curve passing through $x$, and $v\equiv v(x)=(v^{i}(x))_{i\in\{1,2,3\}}$ is a vector field.
@@
We begin by pointing out that, as a consequence of Lemma 7, $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$ are (precise) Markov semigroups, and so the infinitesimal generators $\underline{L}$ and $\overline{L}$ are well-defined.
@@
Here the factor $\frac{1-e^{-2t\rho_{\overline{L}}}}{\rho_{\overline{L}}}$ needs to be understood as $2t$ when $\rho_{\overline{L}}=0$. Following van Handel (2016, Theorem 2.35), taking the limit as $t\rightarrow\infty$ on both sides of (5) yields the following Poincaré inequality
@@
Here it is implicitly assumed that $\rho_{\overline{L}}>0$; We expand on this in Remark 4. In addition, recall that for the Dirichlet form $\overline{\mathcal{E}}(f)$ we have that
@@
Combining (7) with (6), we have that if function $f$ is such that $\overline{L}f=0$, then $\overline{\mathcal{E}}(f)=0$. In turn, this implies that $\text{Var}_{\overline{\mu}}(f)=0$, and so the function $f$ must be constant $\overline{\mu}$-a.e. Moreover, $\mathcal{A}_{0}$ is a core for $\overline{L}$, i.e. it is dense in $\mathcal{D}(\overline{L})$ for the graph norm $\|f\|_{\mathcal{D}(\overline{L})}=\|f\|_{\mathbb{L}^{2}(\overline{\mu})}+\|\overline{L}f\|_{\mathbb{L}^{2}(\overline{\mu})}$. [^10] This norm and its properties are recalled in Appendix C.2. Therefore, equations (3)-(7) established on $\mathcal{A}_{0}$ extend to $\mathcal{D}(\overline{L})$ by approximation. This implies that every $f\in\mathcal{D}(\overline{L})$ such that $\overline{L}f=0$ is constant, and so that $\overline{L}$ is ergodic according to Bakry et al. (2014, Definition 3.1.11). Then, since $\overline{\mu}$ is finite, by Bakry et al. (2014, Proposition 3.1.13) we have that
@@
Theorem 10 is about the chosen representatives $\underline{\mathbf{P}}\in\underline{[\mathbf{P}]}_{\tilde{f}}$ and $\overline{\mathbf{P}}\in\overline{[\mathbf{P}]}_{\tilde{f}}$ : the invariant measures $\underline{\mu},\overline{\mu}$ and generators $\underline{L},\overline{L}$ are associated with these representatives. The proof uses only assumptions (i)-(iii) on the representative generator; Extremality is not used beyond selecting which semigroups are considered. Hence, the same convergence conclusion holds for any representative in $\mathbf{Q}$ (in particular, in the extremal classes) satisfying (i)-(iii). Moreover, if two representatives $\mathbf{P},\mathbf{P}^{\prime}$ lie in the same class $\mathbf{P}]=\left[\mathbf{P}^{\prime}\right]$ and both satisfy the conclusion with invariant measures $\mu,\mu^{\prime}$, then they necessarily satisfy $\int\tilde{f}d\mu=\int\tilde{f}d\mu^{\prime}$, since $P_{t}\tilde{f}=P_{t}^{\prime}\tilde{f}$ for all $t$.
Examples of semigroups for which the conditions (i)-(iii) of Theorem 10 are met are the Ornstein-Uhlenbeck and Laguerre semigroups (whose Markov generators satisfy the Bakry-Émery curvature condition with $\rho>0$; In particular, $\rho=1$ for the former, and $\rho=1/2$ for the latter). In addition, an example of two Markov semigroups having the same invariant measure (as required for (2) to hold) are the Brownian motion with quadratic potential, and the Ornstein-Uhlenbeck semigroup with volatility parameter equals to $\sqrt{2}$. Indeed, for both, the invariant measure is a Gaussian with mean 0, and variance the inverse of the rate of mean reversion.
The following is an important corollary that subsumes the ergodic behavior of the elements of $\mathbf{Q}$ when applied to $\tilde{f}\in B(E)$. We denote by $\underline{\mu}$, $\overline{\mu}$ the invariant measures of the chosen representatives of the extremal classes from Theorem 10.
Corollary 10.1 (Ergodicity of Compact Imprecise Markov Semigroup $\mathbf{Q}$, $E$ Euclidean space or Riemannian manifold).
@@
Let us add a few comments on Corollary 10.1. If $\mu$ is an invariant measure for a (precise) semigroup $\mathbf{P}=(P_{t})_{t\geq 0}$, then by definition $\mu P_{t}=\mu$ for all $t\geq 0$; Equivalently, $\int_{E}P_{t}f\mathrm{d}\mu=\int_{E}f\mathrm{d}\mu$ for every bounded $f$. This is a stationarity statement and holds for every $t$ as soon as the initial law is chosen to be $\mu$. Corollary 10.1, however, is of a different nature: It provides long-term bounds for the conditional expectation $P_{t}\tilde{f}(x)=\mathbb{E}[\tilde{f}(X_{t})\mid X_{0}=x]$ (uniformly in the initial state, and robustly over $\mathbf{Q}$), for a fixed $\tilde{f}$, which is not implied by invariance alone and typically requires mixing/ergodicity assumptions. In particular, in the case $\underline{\mu}=\overline{\mu}=\mu$, the bounds collapse to the same constant,
$\lim_{t\to\infty}\underline{P}_{t}\tilde{f}(x)=\lim_{t\to\infty}\overline{P}_{t}\tilde{f}(x)=\int_{E}\tilde{f}\mathrm{d}\mu,$
so the imprecision about the long-run expectation of $\tilde{f}$ vanishes, even though the short-run values may differ across models in $\mathbf{Q}$. Finally, we emphasize that the limits in Corollary 10.1 are not Cesàro time averages along a single trajectory; Rather, they concern convergence of the semigroup (one-time marginals) as $t\to\infty$. Motivated by (8)–(10), we say that $\mathbf{Q}$ is $\tilde{f}-$lower ergodic if it satisfies (8), $\tilde{f}-$upper ergodic if it satisfies (9), and $\tilde{f}-$ergodic if it satisfies (10).
As we pointed out in Section 2, there are two main differences with respect to the existing literatures on imprecise Markov processes and ergodicity for imprecise probabilities. We are the first to explicitly exploit the geometry of the state space $E$ via the Bakry-Émery curvature condition to derive our results. In addition, in Theorem 10 and Corollary 10.1 we represent the scholar’s ambiguity about the invariant measure and the transition probability via a compact imprecise Markov semigroup $\mathbf{Q}$. We briefly mention that our approach differs slightly from the imprecise ergodic literature (Kozine and Utkin, 2002; Hermans and de Cooman, 2012; Skulj, 2013, 2015; Cerreia-Vioglio et al., 2015; de Cooman et al., 2016; Caprio and Gong, 2023; Caprio and Mukherjee, 2023), where ambiguity around measure $\mu$ is typically represented via a Choquet approach involving a superadditive Choquet capacity.
We conclude this section with a corollary that further informs us on the geometrical properties of $E$ in the ergodic regime of $\mathbf{Q}$, when $E$ is a complete Riemannian manifold. Let us denote by $E^{n}_{K}$ the complete $n$-dimensional simply connected space of constant sectional curvature $K\in\mathbb{R}$, [^11] The Ricci curvature of $E^{n}_{K}$ is constant and equal to (n1)K(n-1)K. and by $B(p,r)$ the ball of radius $r$ at point $p$, defined with respect to the Riemannian distance function. Once again, the values that we study are tied to the choice of the representatives $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$.
Corollary 10.2 (Bishop-Gromov Inequality).
@@
Given our assumptions, we know that $CD(\rho_{\underline{L}},\infty)$ and $CD(\rho_{\overline{L}},\infty)$ both hold. Let us define $\rho_{\star}\coloneqq\min\{\rho_{\underline{L}},\rho_{\overline{L}}\}$, and put $\rho_{\star}=(n-1)K$, so that $K=\rho_{\star}/(n-1)$. The Bakry-Émery curvature assumptions imply that the Ricci tensor at every point of $E$ is bounded from below by $\rho_{\star}$. [^12] This is discussed in Appendix C.1.1. In turn, we have that the Ricci curvature Ric of $E$ is lower bounded by (n1)K(n-1)K, that is, $\text{Ric}\geq(n-1)K$. The result then follows from Bishop (1963). ∎
@@
4 Ergodicity in the General Case – Compact $\mathbf{Q}$
In this section, our goal is to extend Theorem 10 and Corollary 10.1 to the case where $(E,\mathcal{F})$ is an arbitrary good measurable space, while retaining the assumptions that $\mathbf{Q}$ is compact and $\tilde{f}$ is fixed. To do so, we need to introduce the concept of an Imprecise Diffusion Markov Tuple, the imprecise version of a diffusion Markov triple; The latter is recalled in Appendix C.2. Similarly to Section 3, in the following definition, theorem, and corollary we have that $\underline{\mathbf{P}}\equiv\underline{\mathbf{P}}_{\tilde{f}}$, a fixed chosen representative from $\underline{[\mathbf{P}]}_{\tilde{f}}$, and similarly for $\overline{\mathbf{P}}$.